OpenStudy (anonymous):

If A = (0, 0) and B = (6, 3), what is the length of ? A. 7.73 units B. 6.24 units C. 5.20 units D. 6.71 units

OpenStudy (ribhu):

use distance formula.

OpenStudy (anonymous):

i tried and its not working

OpenStudy (anonymous):

i used (x2-x1)+ (y2-y1) squared

OpenStudy (ribhu):

it is \[\sqrt{(X{2}-X{1})^{2}+ (Y{2}-Y{1})^{2}}\]

OpenStudy (ribhu):

plug in the values you will get the ans for sure @mikebyy

OpenStudy (anonymous):

(6-0)2+(3-0)2 12+9=21 squard is 4.5

OpenStudy (ribhu):

no its (6-0)^2 = 36 (3-0)^2 = 9 (36+9)^(1/2) = 45^(1/2)

OpenStudy (anonymous):

sqrt(45)